# Feature Representation Transferring to Lightweight Models via Perception Coherence

### v1

@@ 3.2 Practical implementation with a differentiable ranking operation
To implement our method, we use a simple Monte-Carlo sampling, processing by mini-batch, like any standard machine learning algorithm. Consider a training batch :=$\mathcal{B}:=\{x_{i}\}_{i=1}^{B}$. For each $f\in\{f_{1},f_{2}\}$ endowed with the metrics $d\in\{d_{f_{1}},d_{f_{2}}\}$, we first compute the pair-wise dissimilarity $d(x_{i},x_{j})$ for all $i\neq j$. Let $d_{ij}$ denote $d(x_{i},x_{j})$ for brevity. Let us denote $D^{f}_{i}(\mathcal{B}):=(d_{ij})_{j=1}^{B}$. Let $\mathcal{R}^{f}_{i}(\mathcal{B})$ be the ranking of $D^{f}_{i}(\mathcal{B})$. That is, the $j^{th}$ component of $\mathcal{R}^{f}_{i}(\mathcal{B})$ is the rank of $d_{ij}$ in $D^{f}_{i}(\mathcal{B})$. Notice that $\mathcal{R}^{f}_{i}(\mathcal{B})$ is a permutation of $\{1,2\cdots,B\}$ (so, $\mathcal{R}^{f}_{i}(\mathcal{B})\in\mathbb{R}^{B}$). Based on Remark 2, intuitively, with sufficiently large $B$ [^1] Through our qualitative observation, for $B$ from $\sim 10$, results start to stabilize. See Appendix E.4 for results and discussions., $\frac{\mathcal{R}_{i}^{f_{1}}(\mathcal{B})}{B}\approx\left(F_{1}(x_{i},x_{j})\right)_{j=1}^{B}$ (similarly for $f_{2}$). Hence, to empirically maximize the perception coherence level (Eq. (3.2)), we can minimize
@@
We use here the squared Euclidean norm for its simplicity, but other metrics can be used to compute the difference between $\mathcal{R}_{i}^{f_{1}}(\mathcal{B})$ and $\mathcal{R}_{i}^{f_{2}}(\mathcal{B})$. Besides, notice that each $i\in\{1,2\cdots,B\}$ corresponds to $x_{i}$ being the reference point. Hence, averaging over $i$ in the batch allows for maximizing the global expected coherence level (the expectation is taken over different reference points, see Definition 4 for recalling). However, the ranking operation is not differentiable. Hence, we propose a simple soft ranking operation based on sigmoid function. Assuming that there is no ties, i.e., $d_{ij}\neq d_{ik}$, for all $j\neq k$, that generally holds in practice. In this case, we notice that the rank of $d_{ij}$ in $D_{i}(\mathcal{B})$ can be computed as $r(d_{ij})=\sum_{k=1}^{B}\mathds{1}_{\{d_{ik}\leq d_{ij}\}}$. Hence, we can soften the ranking operation as follows
$\widetilde{r}(d_{ij})=\sum_{k=1}^{B}\Lambda\left(\frac{d_{ij}-d_{ik}}{\tau}\right)\ ,$ (3.3)
where $\Lambda$ is the sigmoid function and $\tau<1$ is the temperature playing the role of scaling. The above function approaches the step function as $\tau$ tends to 00. Hence, this allows us to conveniently approximate the ranking operation. Let ~$\widetilde{\mathcal{R}}_{i}(\mathcal{B})$ be the corresponding ranking obtained with this softened version. So, our loss function can be written as,
$\mathcal{L}_{\text{ours}}(f_{1},f_{2};\mathcal{B}):=\frac{1}{B^{3}}\sum_{i=1}^{B}\left\|\widetilde{\mathcal{R}}_{i}^{f_{1}}(\mathcal{B})-\widetilde{\mathcal{R}}_{i}^{f_{2}}(\mathcal{B})\right\|^{2}\ .$ (3.4)
3.3 Theoretical insights
Throughout our theoretical analysis, we make the following assumption.
Assumption 1.
@@ 3.3 Theoretical insights
In this section, we first derive theoretical insights about the local perception coherence at a given point $x$ in Section 3.3.1. Next, we also investigate about the global coherence in Section 3.3.2. Finally, we study the stability of the perception coherence around a given local region in Section 3.3.3.
3.3.1 Local perception coherence
Theorems 3.1 and 3.2 provide insight into how well the student model $f_{2}$ mimics the teacher model $f_{1}$, given the local perception coherence level.
Theorem 3.1 (Local rank preservation).
@@ 3.3.1 Local perception coherence
$\mathbb{P}_{X,X^{\prime}}\left(\left|F_{2}(x,X^{\prime})-F_{2}(x,X)\right|\geq\varepsilon_{2}\ \Bigm{|}\ \left|F_{1}(x,X^{\prime})-F_{1}(x,X)\right|\leq\varepsilon_{1}\right)\leq\frac{C_{\varepsilon_{1}}(1-\alpha)}{\varepsilon_{2}-\varepsilon_{1}}\ ,$ (3.5)
@@
The proof is in Appendix B.1. Consider the case where $\alpha$ approaches 1, then the right-hand side (RHS) of Eq. (3.5) approaches 0. Theorem 3.5 shows that, if we pick up two random points $x_{1}$ and $x_{2}$ such that the teacher model $f_{1}$ perceives them to have close dissimilarities w.r.t. a given point $x$ (i.e., their dissimilarities differ by at most $\varepsilon_{1}$), then there is a high chance that it is also the case for the student model $f_{2}$. That is, $f_{2}$ also perceives that their dissimilarities differ by at most $\varepsilon_{1}$. This is because if $\alpha=1$, the RHS is equal to 00 for any $\varepsilon_{2}>\varepsilon_{1}$. While this theorem tells us only about the difference between the dissimilarities (absolute value), Theorem 3.2 also considers the ordering of the dissimilarities.
Theorem 3.2 (Local relative order preservation).
@@
$\mathbb{P}_{X,X^{\prime}}\left(F_{2}(x,X^{\prime})-F_{2}(x,X)\geq\varepsilon\ \big{|}\ F_{1}(x,X^{\prime})\leq F_{1}(x,X)\right)\leq C(1-\alpha)/\varepsilon\ ,$ (3.6)
@@
The proof of this theorem can be found in Appendix B.2. Notice that the RHS of Eq. (3.6) approaches 00 as $\alpha$ approaches $1$ for any $\varepsilon>0$. Hence, in this case, Theorem 3.2 shows that, if we pick up two random points $x_{1}$ and $x_{2}$ such that the teacher model $f_{1}$ perceives $x_{1}$ to be more similar to $x$ than $x_{2}$ (i.e., $F_{1}(x,x_{1})\leq F_{1}(x,x_{2})$), then probability that $f_{2}$ perceives $x_{1}$ to be less similar to $x$ than $x_{2}$ is upper bounded. In the case where $\alpha=1$, this probability is 00, that can be formally stated as follows.
Corollary 3.1.
@@
3.3.2 Global perception coherence
@@
The proof of this theorem can be found in Appendix C.1. Theorem 3.3 shows that, even without fixing the reference point, knowing the global expected coherence guarantees a bound. This result can be interpreted similarly to the local perception case in Theorems 3.1 and 3.2.
3.3.3 Stability of perception coherence around a local region
@@
$\mathbb{P}_{X}\left(\phi_{f_{1},f_{2}}(X)\leq\alpha-\varepsilon\ \Bigm{|}\ X\in\mathcal{A}_{\delta}\right)\leq\frac{O(\delta)}{\varepsilon C_{\mathcal{A}}}\ ,$ (3.7)
@@ 4 Experiments
In this section, we first conduct experiments on $2D$ and $3D$ data for proof-of-concept (Section 4.1). Then, in Section 4.2, we conduct experiments on lightweight settings to evaluate the quality of the learned representation through retrieval tasks. Finally, we demonstrate how our method can boost the performance of student models in classification task in Section 4.3. Without further mention, we use cosine dissimilarity as the dissimilarity metric on neural networks (defined in Appendix E). For brevity, the details for all experiments can also be found in Appendix E.
@@ 4.1 Proof-of-concept with experiments on $\mathbf{2D}$ and $\mathbf{3D}$ toy examples
In this section, we conduct experiments on $2D$ and $3D$ data, without neural networks, to qualitatively demonstrate how our method works. For this, we use a reference data configuration $\{x^{1}_{i}\}_{i=1}^{N}$ (can be seen as the teacher configuration, $N$ is the number of data points). Then, we randomly initialize another configuration $\{x^{2}_{i}\}_{i=1}^{N}$ (student configuration), where each point $x^{2}_{i}$ is associated with a fixed point $x^{1}_{i}$. More formally, for each $i\in\{1,2\cdots N\}$, $f_{1}(i)=x^{1}_{i}$ (fixed) and $f_{2}(i)=x^{2}_{i}$. We then apply our loss function from Eq. (3.4), where the dissimilarity metric is the Euclidean distance. All experiment details are in Appendix E.1. The evolution of the student configurations in different setups are shown in Fig. 4.1 and 4.2. We intentionally set the teacher and student configurations in different scales of magnitude to show the effectiveness of the method (teacher configuration has data points in scale $\sim 1$ whereas student configuration is in scale $\sim 10$). We see that our method results in a very smooth configuration transfer, despite the scale difference. From Fig. 4.1, we remark that, while the exact geometry is not completely preserved, the learned configuration preserves a global structural coherence. The same phenomenon is also observed in the case where teacher and student configurations have different dimensions ($3D\to 2D$) in Fig. 4.2. Another experiment with more clusters is in Appendix E.1. Thus, this qualitatively proves the effectiveness of our method.
Figure 4.1: Transferring process on $2D$ datasets. Top row represents the teacher configuration (fixed), bottom row represents the evolution of the student configuration along the transfer process (from left to right). Each gray line associates $x^{2}_{i}$ with $x^{1}_{i}$ (for the same $i$). We observe that the learned configuration preserves a global structural coherence without preserving completely the geometry.
@@
Figure 4.2: Transferring process from $3D$ to $2D$ data.

### v2

@@ 3.2 Practical implementation with a differentiable ranking operation
To implement our method, we use a simple Monte-Carlo sampling, processing by mini-batch, like any standard machine learning algorithm. Consider a training batch $\mathcal{B}:=\{x_{i}\}_{i=1}^{B}$. For each $f\in\{f_{1},f_{2}\}$ endowed with the metrics $d\in\{d_{f_{1}},d_{f_{2}}\}$, we first compute the pair-wise dissimilarity $d(x_{i},x_{j})$ for all $i,j\in\{1,2,\cdots,B\}$. Let $d_{ij}$ denote $d(x_{i},x_{j})$ for brevity. Let us denote $D^{f}_{i}(\mathcal{B}):=(d_{ij})_{j=1}^{B}$. Let $\mathcal{R}^{f}_{i}(\mathcal{B})$ be the ranking of $D^{f}_{i}(\mathcal{B})$. That is, the $j^{th}$ component of $\mathcal{R}^{f}_{i}(\mathcal{B})$ is the rank of $d_{ij}$ in $D^{f}_{i}(\mathcal{B})$. Notice that $\mathcal{R}^{f}_{i}(\mathcal{B})$ is a permutation of $\{1,2\cdots,B\}$ (so, $\mathcal{R}^{f}_{i}(\mathcal{B})\in\mathbb{R}^{B}$). Based on Remark 3.2, intuitively, with sufficiently large BB [^1] Through our qualitative observation, for $B$ from $\sim 10$, results start to stabilize. See Section 5.4 for results and discussions., for each $f\in\{f_{1},f_{2}\}$, we have
$\frac{\mathcal{R}_{i}^{f_{1}}(\mathcal{B})}{B}\approx\left(F_{1}(x_{i},x_{j})\right)_{j=1}^{B}\ ,\;\frac{\mathcal{R}_{i}^{f_{2}}(\mathcal{B})}{B}\approx\left(F_{2}(x_{i},x_{j})\right)_{j=1}^{B}\ .$
To maximize the perception coherence level (Eq. (2)), we need to somehow minimize the discrepancy between $F_{1}$ and $F_{2}$. Thus, empirically, we can minimize
@@
We use here the squared Euclidean norm for its simplicity, but other metrics can be used to compute the difference between $\mathcal{R}_{i}^{f_{1}}(\mathcal{B})$ and $\mathcal{R}_{i}^{f_{2}}(\mathcal{B})$. Besides, notice that each $i\in\{1,2\cdots,B\}$ corresponds to $x_{i}$ being the reference point. Hence, averaging over $i$ in the batch allows for maximizing the global expected coherence level (the expectation is taken over different reference points, see Definition 3.4 for recalling). However, the ranking operation is not differentiable. Hence, we propose a simple soft ranking operation based on sigmoid function. Assuming that there is no ties, i.e., $d_{ij}\neq d_{ik}$, for all $j\neq k$, that generally holds in practice. In this case, we notice that the rank of $d_{ij}$ in $D_{i}(\mathcal{B})$ can be computed as $r(d_{ij})=\sum_{k=1}^{B}\mathds{1}_{\{d_{ik}\leq d_{ij}\}}$. Hence, we can soften the ranking operation as follows
$\widetilde{r}(d_{ij})=\sum_{k=1}^{B}\Lambda\left(\frac{d_{ij}-d_{ik}}{\tau}\right)\ ,$ (3)
where $\Lambda$ is the sigmoid function and $\tau<1$ is the temperature playing the role of scaling. The above function approaches the step function as $\tau$ tends to 0. Hence, this allows us to conveniently approximate the ranking operation. Let $\widetilde{\mathcal{R}}_{i}(\mathcal{B})$ be the corresponding ranking obtained with this softened version. So, our loss function can be written as,
$\mathcal{L}_{\text{ours}}(f_{1},f_{2};\mathcal{B}):=\frac{1}{B^{3}}\sum_{i=1}^{B}\left\|\widetilde{\mathcal{R}}_{i}^{f_{1}}(\mathcal{B})-\widetilde{\mathcal{R}}_{i}^{f_{2}}(\mathcal{B})\right\|^{2}\ .$ (4)
4 Theoretical aspects
In this section, we first investigate in Section 4.1 the convergence behavior of the estimator of perception coherence level using mini-batches. Then, we derive theoretical insights about the local perception coherence at a given point $x$ in Section 4.2. Next, we also investigate about the global coherence in Section 4.3. Finally, we study the stability of the perception coherence around a given local region in Section 4.4. Throughout our theoretical analysis, we make the following assumptions.
Assumption 4.1 (Measurability of Distribution Function)
For $i=1,2$, the dissimilarity function $d_{i}:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ is measurable, and the marginal distribution $\mathcal{D}$ is such that the function $F_{i}(x,x^{\prime}):=\mathbb{P}_{X\sim\mathcal{D}}\big(d_{i}(x,X)\leq d_{i}(x,x^{\prime})\big)$ is jointly measurable in $(x,x^{\prime})\in\mathcal{X}\times\mathcal{X}$.
Assumption 4.2
@@ 4 Theoretical aspects
4.1 Expected Convergence of the Mini-Batch Estimator
Recall from Eq. (2) that the perception coherence level of $f_{2}$ with respect to (w.r.t.) $f_{1}$ at $x\in\mathcal{X}$ is defined as
$\phi_{f_{1},f_{2}}(x):=1-\mathbb{E}_{X\sim\mathcal{D}}\Big[\big|F_{1}(x,X)-F_{2}(x,X)\big|\Big].$
Averaging over the data distribution yields the global perception coherence,
$\mathbb{E}_{X}\!\left[\phi_{f_{1},f_{2}}(X)\right]=1-\mathbb{E}_{X_{1},X_{2}\sim\mathcal{D}}\Big[\big|F_{1}(X_{1},X_{2})-F_{2}(X_{1},X_{2})\big|\Big].$
Maximizing the global perception coherence is equivalent to minimizing the following complementary quantity, which we call the Difference Coefficient:
$\mathrm{DC}:=1-\mathbb{E}_{X}[\phi_{f_{1},f_{2}}(X)]=\mathbb{E}_{X_{1},X_{2}}\Big[\big|F_{1}(X_{1},X_{2})-F_{2}(X_{1},X_{2})\big|\Big].$
Intuitively, $\mathrm{DC}$ measures the discrepancy between the rankings induced by $f_{1}$ and $f_{2}$ over the data distribution (i.e., discrepancy between $F_{1}$ and $F_{2}$).
Mini-batch approximation.
In practice, computing $\mathrm{DC}$ exactly is infeasible, as it requires access to the full distribution (or the entire dataset). Instead, we approximate it using i.i.d. samples drawn in mini-batches. The corresponding empirical estimator is defined as follows.
Definition 4.1 (Empirical Estimator)
Let $\mathcal{B}=\{X_{1},\ldots,X_{B}\}$ be a mini-batch of $B$ samples drawn i.i.d. from $\mathcal{D}$. For each $i\in\{1,2\}$, define the empirical distribution function
$\widehat{F}_{i,B}(x,x^{\prime}):=\frac{1}{B}\sum_{k=1}^{B}\mathds{1}\big\{d_{i}(x,X_{k})\leq d_{i}(x,x^{\prime})\big\}.$
Then the empirical difference coefficient is
$\widehat{\mathrm{DC}}_{B}:=\frac{1}{B^{2}}\sum_{i=1}^{B}\sum_{j=1}^{B}\big|\widehat{F}_{1,B}(X_{i},X_{j})-\widehat{F}_{2,B}(X_{i},X_{j})\big|.$
This estimator depends only on the samples in the mini-batch and provides a tractable approximation to $\mathrm{DC}$. Before analyzing its convergence properties, we first verify that the inner terms are unbiased estimators of the population functions.
Proposition 4.1 (Unbiasedness)
For each fixed $i\in\{1,2\}$ and any pair $(x,x^{\prime})\in\mathcal{X}^{2}$, we have
$\mathbb{E}_{\mathcal{B}}\!\left[\widehat{F}_{i,B}(x,x^{\prime})\right]=F_{i}(x,x^{\prime})\,.$
In particular, $\widehat{F}_{i,B}$ is an unbiased estimator of $F_{i}$.
Proof 4.1
By linearity of expectation,
$\begin{split}\mathbb{E}\!\left[\widehat{F}_{i,B}(x,x^{\prime})\right]&=\frac{1}{B}\sum_{k=1}^{B}\mathbb{E}_{X_{k}\sim\mathcal{D}}\!\left[\mathds{1}\{d_{i}(x,X_{k})\leq d_{i}(x,x^{\prime})\}\right]=\mathbb{E}_{X\sim\mathcal{D}}\!\left[\mathds{1}\{d_{i}(x,X)\leq d_{i}(x,x^{\prime})\}\right]\\ &=\mathbb{P}_{X\sim\mathcal{D}}\!\left(d_{i}(x,X)\leq d_{i}(x,x^{\prime})\right)=F_{i}(x,x^{\prime})\,.\end{split}$
Convergence of the mini-batch estimator.
We now establish that $\widehat{\mathrm{DC}}_{B}$ converges to the true $\mathrm{DC}$ at the standard Monte Carlo rate.
Theorem 4.1 (Expected Convergence Rate)
Let $\widehat{\mathrm{DC}}_{B}$ be the empirical estimator of $\mathrm{DC}$ based on a batch of size $B$. Then there exists a constant $C>0$, independent of $B$, such that
$\mathbb{E}_{\mathcal{B}}\!\left[\,\big|\widehat{\mathrm{DC}}_{B}-\mathrm{DC}\big|\,\right]\leq\frac{C}{\sqrt{B}}\,.$
Consequently, $\widehat{\mathrm{DC}}_{B}\to\mathrm{DC}$ in expectation at the rate $O(1/\sqrt{B})$ as $B\to\infty$.
Proof 4.2
The detailed proof is provided in Appendix A.
Discussion.
Theorem 4.1 ensures that the estimation error decays proportionally to $1/\sqrt{B}$. In other words, enlarging the mini-batch size systematically reduces the variance of the estimator, thereby improving the accuracy of the approximation to $\mathrm{DC}$. This theoretical guarantee is strongly supported by our experiments, that will be shown in Section 5.4 (Fig. 6 and Table 4). More precisely, we shall show that the empirical estimation error exhibits the predicted scaling behavior with respect to $B$.
4.2 Local perception coherence
Theorems 4.2 and 4.3 provide insight into how well the student model $f_{2}$ mimics the teacher model $f_{1}$, given the local perception coherence level.
Theorem 4.2 (Local rank preservation)
@@ 4.2 Local perception coherence
$\mathbb{P}_{X,X^{\prime}}\left(\left|F_{2}(x,X^{\prime})-F_{2}(x,X)\right|\geq\varepsilon_{2}\ \Bigm|\ \left|F_{1}(x,X^{\prime})-F_{1}(x,X)\right|\leq\varepsilon_{1}\right)\leq\frac{C_{\varepsilon_{1}}(1-\alpha)}{\varepsilon_{2}-\varepsilon_{1}}\ ,$ (5)
@@
The proof is in Appendix B.1. Consider the case where $\alpha$ approaches 1, then the right-hand side (RHS) of Eq. (5) approaches 0. Theorem 5 shows that, if we pick up two random points $x_{1}$ and $x_{2}$ such that the teacher model $f_{1}$ perceives them to have close dissimilarities w.r.t. a given point $x$ (i.e., their dissimilarities differ by at most $\varepsilon_{1}$), then there is a high chance that it is also the case for the student model $f_{2}$. That is, $f_{2}$ also perceives that their dissimilarities differ by at most $\varepsilon_{1}$. This is because if $\alpha=1$, the RHS is equal to 0 for any $\varepsilon_{2}>\varepsilon_{1}$. While this theorem tells us only about the difference between the dissimilarities (absolute value), Theorem 4.3 also considers the ordering of the dissimilarities.
Theorem 4.3 (Local relative order preservation)
@@
$\mathbb{P}_{X,X^{\prime}}\left(F_{2}(x,X^{\prime})-F_{2}(x,X)\geq\varepsilon\ \big|\ F_{1}(x,X^{\prime})\leq F_{1}(x,X)\right)\leq C(1-\alpha)/\varepsilon\ ,$ (6)
@@
The proof of this theorem can be found in Appendix B.2. Notice that the RHS of Eq. (6) approaches 0 as $\alpha$ approaches $1$ for any $\varepsilon>0$. Hence, in this case, Theorem 4.3 shows that, if we pick up two random points $x_{1}$ and $x_{2}$ such that the teacher model $f_{1}$ perceives $x_{1}$ to be more similar to $x$ than $x_{2}$ (i.e., $F_{1}(x,x_{1})\leq F_{1}(x,x_{2})$), then probability that $f_{2}$ perceives $x_{1}$ to be less similar to $x$ than $x_{2}$ is upper bounded. In the case where $\alpha=1$, this probability is 0, that can be formally stated as follows.
Corollary 4.1
@@
4.3 Global perception coherence
@@
The proof of this theorem can be found in Appendix C.1. Theorem 4.4 shows that, even without fixing the reference point, knowing the global expected coherence guarantees a bound. This result can be interpreted similarly to the local perception case in Theorems 4.2 and 4.3.
4.4 Stability of perception coherence around a local region
@@
$\mathbb{P}_{X}\left(\phi_{f_{1},f_{2}}(X)\leq\alpha-\varepsilon\ \Bigm|\ X\in\mathcal{A}_{\delta}\right)\leq\frac{O(\delta)}{\varepsilon C_{\mathcal{A}}}\ ,$ (7)
@@ 5 Experiments
In this section, we first conduct experiments on $2D$ and $3D$ data for proof-of-concept (Section 5.1). Then, in Section 5.2, we conduct experiments on lightweight settings to evaluate the quality of the learned representation through retrieval tasks. Finally, we demonstrate how our method can boost the performance of student models in classification task in Section 5.3. We shall also conduct ablation study in Sections 5.4 and 5.5. Without further mention, we use cosine dissimilarity as the dissimilarity metric on neural networks (defined in Appendix E). For brevity, the details for all experiments can also be found in Appendix E.
@@ 5.1 Proof-of-concept with experiments on $\mathbf{2D}$ and $\mathbf{3D}$ toy examples
In this section, we conduct experiments on $2D$ and $3D$ data, without neural networks, to qualitatively demonstrate how our method works. For this, we use a reference data configuration $\{x^{1}_{i}\}_{i=1}^{N}$ (can be seen as the teacher configuration, $N$ is the number of data points). Then, we randomly initialize another configuration $\{x^{2}_{i}\}_{i=1}^{N}$ (student configuration), where each point $x^{2}_{i}$ is associated with a fixed point $x^{1}_{i}$. More formally, for each $i\in\{1,2\cdots N\}$, $f_{1}(i)=x^{1}_{i}$ (fixed) and $f_{2}(i)=x^{2}_{i}$. We then apply our loss function from Eq. (4), where the dissimilarity metric is the Euclidean distance. All experiment details are in Appendix E.1. The evolution of the student configurations in different setups are shown in Fig. 4 and 5. We intentionally set the teacher and student configurations in different scales of magnitude to show the effectiveness of the method (teacher configuration has data points in scale $\sim 1$ whereas student configuration is in scale $\sim 10$). We see that our method results in a very smooth configuration transfer, despite the scale difference. From Fig. 4, we remark that, while the exact geometry is not completely preserved, the learned configuration preserves a global structural coherence. The same phenomenon is also observed in the case where teacher and student configurations have different dimensions ($3D\to 2D$) in Fig. 5. Another experiment with more clusters is in Appendix E.1. Thus, this qualitatively proves the effectiveness of our method.
Figure 4: Transferring process on $2D$ datasets. Top row represents the teacher configuration (fixed), bottom row represents the evolution of the student configuration along the transfer process (from left to right). Each gray line associates $x^{2}_{i}$ with $x^{1}_{i}$ (for the same ii). We observe that the learned configuration preserves a global structural coherence without preserving completely the geometry.
@@
Figure 5: Transferring process from $3D$ to $2D$ data.